OpenStudy (anonymous):

Does the equation represent a direct variation? If so, find the constant of variation. 5x = -3y A. no B. yes; k = -5/3 C. yes; k = 5/3 D. yes; k = 3/5

OpenStudy (anonymous):

Direct variation means the equation is linear.The above equation is linear. And the form is y = kx. y = -5x/3 .So k = -5/3. B. yes; k = -5/3 is the answer.

OpenStudy (anonymous):

kartik is right

OpenStudy (anonymous):

If y can stand alone then the function is linear. So by dividing 5x by -3 y stands alone and k=-5/3